OpenStudy (anonymous):

The diameter of the circle Michael drew was 14.4 cm. What was the approximate circumference of the circle if you use 3.14 for ?

OpenStudy (espex):

Knowing that the circumference of a circle is \(2\pi r\) or \(\pi d\), and pi = 3.14, you are left with a simple multiplication problem.

OpenStudy (espex):

I am not sure what you need here, you have the diameter, which is equal to twice the radius, and you have pi. Fill in the blanks.

OpenStudy (espex):

Your PMQ, "There is a circular pool in the backyard with an approximate area of 113 square inches. What is the radius?" can be solved with the same process. Your area equation is \(A=\pi r^2\) Plug in your A of 113, pi of 3.14 and solve for 'r'.